### v5

# Monte Carlo Neural PDE Solver for Learning PDEs via Probabilistic Representation

Abstract
In scenarios with limited available data, training the function-to-function neural PDE solver in an unsupervised manner is essential. However, the efficiency and accuracy of existing methods are constrained by the properties of numerical algorithms, such as finite difference and pseudo-spectral methods, integrated during the training stage. These methods necessitate careful spatiotemporal discretization to achieve reasonable accuracy, leading to significant computational challenges and inaccurate simulations, particularly in cases with substantial spatiotemporal variations. To address these limitations, we propose the Monte Carlo Neural PDE Solver (MCNP Solver) for training unsupervised neural solvers via the PDEs’ probabilistic representation, which regards macroscopic phenomena as ensembles of random particles. Compared to other unsupervised methods, MCNP Solver naturally inherits the advantages of the Monte Carlo method, which is robust against spatiotemporal variations and can tolerate coarse step size. In simulating the trajectories of particles, we employ Heun’s method for the convection process and calculate the expectation via the probability density function of neighbouring grid points during the diffusion process. These techniques enhance accuracy and circumvent the computational issues associated with Monte Carlo sampling. Our numerical experiments on convection-diffusion, Allen-Cahn, and Navier-Stokes equations demonstrate significant improvements in accuracy and efficiency compared to other unsupervised baselines.
Index Terms:
Neural PDE solver, Monte Carlo method, Feynman-Kac formula, AI for PDE.
1 Introduction
Solving partial differential equations (PDEs) is essential for understanding and modeling various physical phenomena, including fluid dynamics [2], heat transfer [72], and quantum mechanics [26]. Traditional numerical methods for PDEs, such as finite difference, finite element, and spectral methods, have achieved remarkable successes and are widely used across various industries [38]. However, they still have certain limitations. For instance, these methods often require precise spatiotemporal discretization, which can be computationally expensive, particularly in scenarios with high spatiotemporal variability [13]. Additionally, they may struggle with complex geometries and multiscale problems, where the need for specialized mesh designing and treatment can complicate the simulation process and increase user difficulty [27, 94].
Recently, deep learning (DL) methods have emerged as a promising approach to addressing scientific computation challenges for PDE solving, offering a new perspective beyond traditional numerical techniques [32]. By harnessing the representation power of deep neural networks, DL-based techniques have successfully overcome several limitations inherent in classical numerical strategies. For instance, physics-informed neural networks (PINNs) [68, 69] and deep energy methods (DEM) [75, 58] respectively utilize the strong or weak form of PDEs to construct loss functions, providing flexibility in handling complex boundary problems without the need for mesh discretization. Additionally, DL-based reduced-order modeling (ROM) [10, 16] improves the accuracy and efficiency of traditional ROM techniques, particularly in tackling nonlinear and multiscale problems. However, these methods are typically designed for PDEs with fixed initial or forcing fields, necessitating retraining of the neural networks when handling new initial fields.
Beyond these achievements, the function-to-function neural PDE solvers have risen as another new paradigm for simulating physical systems, which leverage neural networks as surrogate models to approximate the solutions of a family of PDEs [76, 87, 48, 43]. Along this direction, several studies have proposed diverse network architectures for neural PDE solvers [76, 42, 48, 7, 98]. These solvers can be trained using supervised [76, 48, 42] or unsupervised approaches [87, 90, 45], employing pre-generated data or PDE information to construct training targets, respectively. The unsupervised training approach is essential for DL-based PDE solvers, particularly in scenarios with limited available or high-quality data. To address this issue, some studies [87, 80, 45] borrow techniques from classical numerical solvers to construct training targets. For instance, the MAC grid physics-constrained network and its 3D extension [87, 88], low-rank decomposition network (LordNet) [80] and physics-informed neural operator (PINO) [45] integrate finite difference or pseudo-spectral methods with neural networks during the training stage. However, these traditional Eulerian methods require fine meshes or step sizes for stable simulations. Therefore, the performance and efficiency of corresponding neural PDE solvers are also limited by time and space discretization, particularly when handling high spatiotemporal variations. Furthermore, these methods often necessitate additional loss functions to enforce boundary conditions, introducing extra complexity in hyperparameter tuning and computational demands.
To this end, we introduce the Monte Carlo Neural PDE Solver (MCNP Solver), a novel approach for training neural solvers from a probabilistic perspective, which views macroscopic phenomena as ensembles of random movements of microscopic particles [96]. For a PDE system with probabilistic representation, we construct the loss function of the MCNP Solver between two sequential PDE fields $u_{t}$ and $u_{t+\Delta t}$ via the relationship derived from the Monte Carlo approximation. To ensure the efficiency and accuracy of the MCNP Solver, we develop several techniques when combining the Monte Carlo approximation with the training of the deep neural network. Specifically, in simulating the corresponding stochastic difference equations (SDE), we use the Heun’s method to obtain a more accurate result when handling the convection process. During the diffusion process, we approximate the mathematical expectation via the probability density function (PDF) of neighbouring grid points to eliminate sampling a large number of particles in Monte Carlo methods. Compared to other unsupervised neural solvers, such as LordNet [80] and PINO [45], the MCNP Solver naturally inherits the advantages of Monte Carlo methods. It can tolerate coarse step size [17, 54], thereby reducing training costs and accumulated errors arising from temporal discretization. Additionally, it can efficiently handle high-frequency spatial fields due to its derivative-free property [4, 50]. Moreover, the boundary conditions are automatically encoded into the stochastic process of particles [4, 50], eliminating the need to introduce extra loss terms to satisfy such constraints.
In summary, we make the following contributions:
1. We propose the MCNP Solver, an innovative unsupervised approach for training neural solvers that can be applied to PDE systems with probabilistic representation. We also devise several strategies to boost performance and efficiency during the convection and diffusion processes in SDE simulations.
2. Our experiments involving the convection-diffusion, Allen-Cahn, and Navier-Stokes equations demonstrate significant improvements in accuracy and efficiency over other unsupervised neural solvers, particularly for simulation tasks with complex spatiotemporal variations and coarse step sizes. Furthermore, we conduct experiments to solve 2D fractional diffusion on a disk, extending the MCNP Solver to mesh-free and fractional Laplacian scenarios.
3. Beyond comparisons with unsupervised learning methods, we comprehensively compare MCNP Solver with other widely-used PDE solvers, including the Eulerian solver, Monte Carlo methods, and the supervised training approach. We delve into a detailed discussion on each method’s strengths, weaknesses, and application scopes.
The structure of this paper is as follows: Section 2 introduces related works on two types of neural PDE solvers. Section 3 provides a detailed overview of the probabilistic representation of the PDEs and the proposed MCNP Solver. Sections 4 and 5 present the experimental results of the MCNP Solver in comparison with unsupervised learning methods and other popular PDE solvers, respectively. Finally, Section 6 summarizes our method, discusses its limitations and outlines potential future work.
2 Related Work
In this section, we introduce two primary categories of neural PDE solvers as follows. The first category is designed to learn a location-to-value mapping for a specific PDE, such as PINN [68]. The second category is designed to learn a function-to-function mapping for a family of PDEs, such as Fourier neural operator (FNO) and DeepONet [43, 48]. In this paper, we target the second task and aim to learn neural PDE solvers between functional spaces that can generalize to different PDE conditions over a distribution.
2.1 The Location-to-Value Neural PDE Solver
The location-to-value PDE solver utilizes the neural network to approximate the solution for a specific PDE with fixed initial and boundary conditions. The input of the neural network is the location and time $(\boldsymbol{x},t)$, and the output is the corresponding solution $u(\boldsymbol{x},t)$. To this end, PINNs have been proposed to construct the loss function using the equation and boundary residuals via the automatic differentiation regime. They are widely employed for solving forward or inverse problems [68, 10, 32, 101]. Recently, PINNs have made significant progress in addressing scientific problems based on PDEs, such as Navier-Stokes equations [69, 52], Schrödinger equations [25, 40], Allen-Cahn equations [53, 31]. Beyond the original PINNs, several methods have been proposed further to enhance the solver accuracy and efficiency of PINNs. For instance, by incorporating more advanced optimization algorithms, the convergence speed of PINNs can be significantly accelerated [30, 49]. Also, several improved architectures can boost PINNs’ capacity to fit high-frequency signals [47, 8]. Some works have sought to improve the performance of PINNs through the design of novel activation functions [18], adaptive collocation strategies [3] and adaptive hyperparameter selection methods [95]. Instead of PINNs, some works utilize the probabilistic representation to train neural networks [23, 20, 97], which can efficiently handle high-dimensional or fractional PDEs [22, 71, 20, 70, 59]. Furthermore, some studies design loss functions based on other numerical methods, such as the finite volume method [5], finite element method [55, 61], and energy-based method [75, 58, 93]. Notably, the aforementioned location-to-value methods require retraining neural networks when encountering a PDE with new initial conditions, which can be time-consuming. In this paper, we aim to learn a function-to-function PDE solver that can generalize over a distribution.
2.2 The Function-to-Function Neural PDE Solver
This kind of neural PDE solver has been proposed to learn mappings between functional spaces, such as mapping a PDE’s initial condition to its solution [48]. Works like DeepONet [48] and its variants [78, 86, 39] encode the initial conditions and queried locations using branch and trunk networks, respectively. Additionally, Fourier Neural Operator (FNO) [43] and its variants [44, 67, 84] explore learning the operator in Fourier space, an efficient approach for handling different frequency components. Several studies have employed graph neural networks [85, 76, 42, 7] or transformers [9, 41] as the backbone models of neural solvers to adapt to complex geometries. However, these methods require the supervision of ground truth data generated via accurate numerical solvers, which can be time-consuming in general. To this end, some studies aim to train the neural PDE solvers without the supervision of data [87, 90, 80, 45]. For example, [90] proposed PI-DeepONets, which utilize the PDE residuals to train DeepONets in an unsupervised way. Similarly, [28] proposed Meta-Auto-Decoder, a meta-learning approach to learn families of PDEs in the unsupervised regime. Furthermore, the MAC grid physics-constrained network [87], LordNet [80], and PINO [45] borrow techniques from traditional Eulerian methods, such as the finite difference and pseudo-spectral methods, and utilize the corresponding residuals as training loss, respectively. In addition to the above approaches, several advanced methods have been proposed to further improve the accuracy of spatiotemporal derivative calculations in unsupervised solver learning methods. For instance, the spline-PINN methods [89, 82] use Hermite spline kernels to interpolate the spatiotemporal field continuously, and the SP-PINN [56] and its extension [57] employ stochastic projection to enhance the precision of spatial gradient computation. Compared to these unsupervised methods, the MCNP Solver integrates physical knowledge via a novel probabilistic perspective, leveraging the strengths of Lagrangian approaches. This approach allows the neural PDE solver to exhibit robustness against spatiotemporal variants and tolerate coarse step size. Also, the boundary conditions and geometric shapes of the PDEs are naturally embedded into the particles’ random walk, eliminating the need for extra constraints in the loss function. Finally, the MCNP Solver can efficiently handle fractional Laplacian operators [34], addressing a gap in the existing unsupervised methods.
3 Methodology
This section introduces the methodology part of the MCNP Solver. Section 3.1 presents the Monte Carlo method and its corresponding theory. Section 3.2 presents the overall framework of the neural Monte Carlo loss, and Section 3.3 provides corresponding details in simulating the convection and diffusion processes in neural Monte Carlo loss.
3.1 Preliminary
Figure 1: Illustration of Feynman-Kac law and the random walks of particles when hitting different boundaries. A: $M$ particles $\{{\boldsymbol{\xi}}_{s}^{m}\}_{m=1}^{M}$ begin at the location $\boldsymbol{x}$, and conduct the random walk according to Eq. 4 from $t=T$ to $t=0$. When $t=0$, the particles query the value at $u_{0}$, and their average value can be considered as the approximation of $u_{T}(x)$ (Eq. 5). B: Monte Carlo methods can naturally encode boundary conditions in the random walk of particles. For periodical/Dirichlet/Neumann boundary conditions, the random walks of particles need to be pulled back/stopped/reflected when hitting the boundaries, respectively.
In this paper, we consider the general convection-diffusion equation defined as follows:
$\frac{\partial u}{\partial t}=\boldsymbol{\beta}[u](\boldsymbol{x},t)\cdot{\nabla u}+\kappa{\Delta u}+f(\boldsymbol{x},t),$ (1)
$u(\boldsymbol{x},0)=u_{0}(\boldsymbol{x}),$
where $\boldsymbol{x}\in\Omega\subset\mathbb{R}^{d}$ and $t$ respectively denote the spatial variable and the temporal variable, $\boldsymbol{\beta}[u](\boldsymbol{x},t)\in\mathbb{R}^{d}$ is a vector-valued mapping from $(u,\boldsymbol{x},t)$ to $\mathbb{R}^{d}$, $\kappa\in\mathbb{R}^{+}$ is the diffusion parameter, and $f(\boldsymbol{x},t)\in\mathbb{R}$ denotes the force term. $\nabla u$ and $\Delta u$ demote the gradient and Laplacian of $u$, respectively. $u_{0}(\boldsymbol{x})$ represents the initial condition. Many well-known PDEs, such as the Allen–Cahn and Navier-Stokes equations, can be viewed as a special form of Eq. 1.
For such PDEs with the form of Eq. 1, the Feynman-Kac formula provides the relationship between the PDEs and the corresponding probabilistic representation [62, 63, 22]. In detail, we can use the time inversion (i.e., $\tilde{u}(\boldsymbol{x},t)=u(\boldsymbol{x},T-t),\tilde{f}(\boldsymbol{x},t)=f(\boldsymbol{x},T-t)$) to the PDE as:
$\frac{\partial\tilde{u}}{\partial t}=-\boldsymbol{\beta}[\tilde{u}](\boldsymbol{x},t)\cdot{\nabla\tilde{u}}-\kappa\Delta\tilde{u}-\tilde{f}(\boldsymbol{x},t),$ (2)
$\tilde{u}(\boldsymbol{x},T)=u_{0}(\boldsymbol{x}).$
Applying the Feynman-Kac formula [51] to the terminal value problem Eq. 2, we have
[…]
where $\tilde{\boldsymbol{\xi}}_{s}\in\mathbb{R}^{d}$ is a random process starting at $\boldsymbol{x}$, and moving from $t=0$ to $t=T$, which satisfies:
$d\tilde{\boldsymbol{\xi}}_{s}=\boldsymbol{\beta}[\tilde{u}](\tilde{\boldsymbol{\xi}}_{s},s)ds+\sqrt{2\kappa}d\boldsymbol{B}_{s},$ (4)
$\tilde{\boldsymbol{\xi}}_{{0}}=\boldsymbol{x},$
where $\boldsymbol{B}_{s}$ is the $d$-dimensional standard Brownian motion. Applying time inversion $t\rightarrow T-t$ to Eq. 3 and letting $\boldsymbol{\xi}$ be the inversion of $\tilde{\boldsymbol{\xi}}$, we have
[…]
We illustrate the diagram of Feynman-Kac law in the 1D case in Fig. 1.A, where the mathematical expectation in Eq. 5 can be naturally approximated via the Monte Carlo sampling. Feynman-Kac formula can automatically encode boundary conditions into the random walk of particles, including the periodical, Dirichlet and Neumann boundary conditions, as discussed in Fig. 1.B. Apart from Eq. 1, some other PDEs can also be handled via the Feynman-Kac formula after specific processing, like wave equations [14] and spatially varying diffusion equations [77].
Based on the Feynman-Kac law, various Monte Carlo methods have been developed to solve PDEs with the form of Eq. 1 [11, 65]. For linear PDEs, Monte Carlo methods can directly simulate the corresponding SDEs and obtain the solution of the targeted PDEs with Eq. 5. For nonlinear PDEs (like Navier-Stokes and Allen-Cahn equations), we cannot simulate the SDEs in Eq. 4 or the mathematical expectation in Eq. 5 directly because the unknown $u$ is required during the simulation. To this end, some advanced numerical algorithms have been developed, such as the random vortex method [66] and the branching diffusion method [24]. Compared to the traditional Eulerian methods, Monte Carlo methods are more adaptable to significant spatiotemporal variations due to their Lagrangian nature, which is widely used for simulating the breaking wave and turbulent problems [29, 35]. Moreover, Monte Carlo methods have been proven to be less restrictive in terms of step size constraints via the analysis of the Courant number [12, 74]. However, the accuracy of Monte Carlo methods is limited by the number of particles, which can introduce severe computational time and memory issues [54]. In the following, we will introduce the MCNP Solver, which can be efficiently trained while inheriting the advantages of the Monte Carlo simulation method.
3.2 Neural Monte Carlo Loss
Given a PDE in the form of Eq. 1 and the distribution of initial conditions $\mathcal{D}_{0}$, the objective of MCNP Solver is to learn a functional mapping $\mathcal{G}_{\theta}$ with parameter $\theta$ that can simulate the subsequent fields for all initial fields $u_{0}\sim\mathcal{D}_{0}$ at time $t\in[0,T]$. The inputs and outputs of $\mathcal{G}_{\theta}$ are given as:
$\mathcal{G}_{\theta}:\mathcal{D}_{0}\times[0,T]$ $\to\mathcal{D}_{[0,T]},$ (6)
$\left(u_{0},t\right)$ $\mapsto u_{t},$
where $\mathcal{D}_{[0,T]}$ denotes the joint distribution of the field after $t=0$. In this paper, we are interested in the evolution of PDEs at fixed coordinate system $\{\boldsymbol{x}_{p}\}_{p=1}^{P}\in\Omega$, which aligns with the settings in [43, 7]. Consequently, the inputs and outputs of the solver $\mathcal{G}_{\theta}$ are solution values at $P$ grid points, i.e., each $u_{t}$ is represented by a $P$-dimensional vector $[u_{t}(\boldsymbol{x}_{1}),\cdots,u_{t}(\boldsymbol{x}_{P})]$.
In practice, we use neural networks [43] to construct the mapping $\mathcal{G}_{\theta}$ and the trained model served as the MCNP Solver. Unlike other supervised learning algorithms [43, 48, 7], MCNP Solver is trained in an unsupervised way, i.e., only utilizing the physics information provided by PDEs. To this end, we consider training the solver via the relationship between $u_{t}$ and $u_{t+\Delta t}$ (where $0\leq t<t+\Delta t\leq T$) derived by the aforementioned probabilistic representation. Considering Eq. 5, the optimal parameter $\theta^{*}$ for the neural PDE solver $\mathcal{G}_{\theta}$ should satisfy the following equation at each $\boldsymbol{x}_{p}$ after optimization:
$\mathcal{G}_{\theta^{*}}(u_{0},t+\Delta t)(\boldsymbol{x}_{p})$ (7)
[…]
where $\boldsymbol{\xi}_{p,s}(s\in[t,t+\Delta t]$) is the inverse version of stochastic process in Eq. 4 as follows:
$d\boldsymbol{\xi}_{p,s}=-\boldsymbol{\beta}[{u}](\boldsymbol{\xi}_{p,s},s)ds-\sqrt{2\kappa}d\boldsymbol{B}_{s},$ (8)
$\boldsymbol{\xi}_{p,{t+\Delta t}}=\boldsymbol{x}_{p}.$
To search for $\theta^{*}$, we convert Eq.  7 to the optimization objective, and thus, the neural Monte Carlo loss can be written as follows:
$\mathcal{L}_{\operatorname{MCNP}}(\mathcal{G}_{\theta}|u_{0})=\sum_{t=0}^{T-\Delta t}\sum_{p=1}^{P}\Bigg{\|}\mathcal{G}_{\theta}(u_{0},t+\Delta t)(\boldsymbol{x}_{p})-$ (9)
[…]
where $\mathcal{G}_{\theta}(u_{0},0)$ is directly set to be $u_{0}$ in experiments. During the training stage, we uniformly sample $B$ initial states $u_{0}$ from $\mathcal{D}_{0}$ per epoch to ensure the generalizability.
Figure 2: Illustration of neural Monte Carlo loss. We construct the training loss via the relationship between ${u}_{t}$ and $u_{t+\Delta t}$ given by the Feynman-Kac law. Given a grid point at $\boldsymbol{x}_{p}$, we split the random walk in Eq. 10 as the convection and diffusion parts. For the convection process, we utilize the Heun’s method to simulate the particle moving from time $t+\Delta t$ to $t$ driven by the drift term $\boldsymbol{\beta}$. For the diffusion process, we calculate the mathematical expectation in Eq. 7 through the probability density function (PDF) of neighbouring grid points to eliminate sampling using Monte Carlo methods, where $p_{t}(\boldsymbol{\xi}_{p,t}^{d,i})$ denotes the transition probability for the particle moving from $\boldsymbol{\xi}_{p,t}^{c}$ to $\boldsymbol{\xi}_{p,t}^{d,i}$ driven by the diffusion effect. Please note that we omit external forcing $f$ in the figure for simplification.
3.3 The Simulation of SDE
When calculating the loss function defined in Eq. 9, a crucial step is to approximate the expectation over random samples $\boldsymbol{\xi}_{p}$ via simulating the SDE in Eq. 8. The Classical Euler–Maruyama method [33] has been adopted to approximate corresponding SDEs [22, 59], i.e.,
${\boldsymbol{\xi}}_{p,t}^{m}$ $={\boldsymbol{\xi}}_{p,t+\Delta t}+\underbrace{\boldsymbol{\beta}[u](\boldsymbol{\xi}_{p,t+\Delta t},t+\Delta t)\Delta t}_{\text{convection}}+\underbrace{\sqrt{2\kappa\Delta t}\boldsymbol{b}^{m}}_{\text{diffusion}},$ (10)
$\boldsymbol{b}^{m}$ $\sim\mathcal{N}(\boldsymbol{0},\boldsymbol{I}),\ {\boldsymbol{\xi}}_{p,t+\Delta t}=\boldsymbol{x}_{p},$
where the physical meaning of $\boldsymbol{\beta}[u](\boldsymbol{\xi}_{p,t+\Delta t},t+\Delta t)\Delta t$ and $\sqrt{2\kappa\Delta t}\boldsymbol{b}^{m}$ denote the convection and diffusion processes, respectively. After sampling $M$ particles $\{\boldsymbol{\xi}^{m}_{p,t}\}_{m=1}^{M}$, the training target in Eq. 7 can be approximated via the Monte Carlo method, i.e.,
$\hat{u}_{t+\Delta t}(\boldsymbol{x}_{p})\approx\frac{1}{M}\sum_{m=1}^{M}\left[u_{t}(\boldsymbol{\xi}_{p,t}^{m})+f(\boldsymbol{\xi}_{p,t+\Delta t},t+\Delta t)\Delta t\right].$ (11)
The selection of the simulation algorithm significantly impacts the efficiency of the training process and the accuracy of the trained solver. This necessitates a careful design approach that considers both the approximation error and the training efficiency, such as computational time and memory. On the one hand, to reduce the discretization error when dealing with the convection term, we consider a higher-order approximation scheme. On the other hand, sampling a large number of particles to calculate the diffusion term will introduce severe computational issues, especially when the diffusive rate $\kappa$ is high. As a result, we explore a method that avoids sampling of the particles during the diffusion process.
In the subsequent sections, we will introduce two methods, including: 1). Utilize Heun’s method to obtain a more accurate approximation of the convection process; 2). Calculate the mathematical expectation via the probability density function (PDF) of neighbouring grid points when handling the diffusion process to eliminate the sampling in Monte Carlo methods, as shown in Fig. 2.
3.3.1 The Simulation of Convection Process
[…]
$\boldsymbol{\beta}_{p,t}^{c,1}$ $=\boldsymbol{\beta}[u]\left(\boldsymbol{\xi}_{p,t+\Delta t},t+\Delta t\right)\Delta t;$ (12)
$\boldsymbol{\beta}_{p,t}^{c,2}$ $=\boldsymbol{\beta}[u]\left(\boldsymbol{\xi}_{p,t+\Delta t}+\boldsymbol{\beta}_{p,t}^{c,1},t\right)\Delta t;$
$\boldsymbol{\xi}_{p,t}^{c}$ $=\boldsymbol{\xi}_{p,t+\Delta t}+\frac{1}{2}\left(\boldsymbol{\beta}_{p,t}^{c,1}+\boldsymbol{\beta}_{p,t}^{c,2}\right).$
For the cases in which the drift term $\boldsymbol{\beta}$ depends on solution $u$, we utilize the output of MCNP Solver to approximate $\boldsymbol{\beta}[u]$ accordingly.
3.3.2 The Simulation of Diffusion Process
Unlike the Euler-Maruyama method, which samples $M$ particles to simulate the diffusion process, we calculate the mathematical expectation via the PDF of neighbouring grid points to replace the sampling in Monte Carlo methods. Given $\boldsymbol{\xi}_{p,t}^{c}$, which represents the location of particles after the convection process (Eq. 12), we first find its neighbourhood $N(\boldsymbol{\xi}_{p,t}^{c},r)$ from the coordinates as follows:
$N(\boldsymbol{\xi}_{p,t}^{c},r)\triangleq\{\boldsymbol{\xi}_{p,t}^{d,i}:\|\boldsymbol{\xi}_{p,t}^{d,i}-\boldsymbol{\xi}_{p,t}^{c}\|_{2}\leq r,\boldsymbol{\xi}_{p,t}^{d,i}\in\{\boldsymbol{x}_{p}\}_{p=1}^{P}\},$ (13)
where $r>0$ represents the radius of the neighbourhood. Considering the transition probability for $\boldsymbol{\xi}_{p,t}^{c}$ moving to its neighbour $\boldsymbol{\xi}_{p,t}^{d,i}\in N(\boldsymbol{\xi}_{p,t}^{c},r)$, we can calculate the analytical form of the corresponding PDF as follows:
$p_{t}(\boldsymbol{\xi}_{p,t}^{d,i})=\frac{1}{({2\pi\sigma^{2}})^{d/2}}\exp\left({-\frac{\|\boldsymbol{\xi}_{p,t}^{d,i}-\boldsymbol{\xi}_{p,t}^{c}\|_{2}^{2}}{2\sigma^{d}}}\right),$ (14)
where $\sigma=\sqrt{2\kappa\Delta t}$ denotes the standard deviation of the Brownian motion in Eq. 8. Please note in Eq. 14, we assume that the particles will not hit the boundary by default. When the particle $\boldsymbol{\xi}_{p,t}^{c}$ has the possibility of hitting the boundary, we need to modify the PDF in Eq. 14 to the corresponding form. For instance, the PDF corresponds to the Brownian motion with absorption/reflection when we handle the Dirichlet/Neumann boundary conditions, respectively. We use the grid points inside the neighbourhood to approximate the mathematical expectation in Eq. 7 as follows:
$\mathbb{E}_{\boldsymbol{\xi}_{p}}\left[\mathcal{G}_{\theta}(u_{0},t)(\boldsymbol{\xi}_{p,t})\right]$ (15)
$\approx$ $\sum_{\boldsymbol{\xi}_{p,t}^{d,i}\in N(\boldsymbol{\xi}_{p,t}^{c},r)}\mathcal{G}_{\theta}(u_{0},t)(\boldsymbol{\xi}_{p,t}^{d,i})\cdot p_{t}(\boldsymbol{\xi}_{p,t}^{d,i})\cdot\delta,$
where we use the density value at point $\boldsymbol{\xi}_{p,t}^{d,i}$ to approximate the density of the cells centered on $\boldsymbol{\xi}_{p,t}^{d,i}$ (the red box in Fig. 3.A) and $\delta$ denotes the volume of each cell. We utilize a similar way to calculate the integration of the external forcing $f$ in Eq. 7 as follows:
[…]
$\approx$ $\frac{\Delta t}{2}\left[f(\boldsymbol{\xi}_{p,t+\Delta t},t+\Delta t)+\sum_{\boldsymbol{\xi}_{p,t}^{d,i}\in N(\boldsymbol{\xi}_{p,t}^{c},r)}f(\boldsymbol{\xi}_{p,t}^{d,i},t)p_{t}(\boldsymbol{\xi}_{p,t}^{d,i})\right].$
Figure 3: The Choice of Neighbourhood Radius $r$ A: We first choose the smallest possible value of $r$ that satisfies Eq. 17. If we have $\sum_{i}p_{t}(\boldsymbol{\xi}_{p,t}^{d,i})\delta\approx 1$, using Eq. 15 directly to approximate the corresponding mathematical expectation can reduce computational cost while ensuring accuracy. $\delta$ denotes the volume of each cell (as shown in the red box) in the coordinate system. B: When the grid size is close to the radius $r$, we may encounter the scenario that $\sum_{i}p_{t}(\boldsymbol{\xi}_{p,t}^{d,i})\delta\ll 1$. To address this issue, we interpolate the coordinate system to a high-resolution one to satisfy the normalization condition in Eq. 17. $\bar{\delta}$ denotes the volume of each cell in the high-resolution coordinate system.
3.3.3 The Choice of Neighbourhood Radius $r$
In Eq. 15, we utilize the grid points located in the neighbourhood $N(\boldsymbol{\xi}_{p,t}^{c},r)$ to approximate the mathematical expectation. To ensure the approximation error is under control, we hope the accumulated probability of the points in the range will approach 1, i.e., given a sufficiently small $\epsilon$, we hope the radius $r$ satisfies the following conditions:
$r=\arg\min{\gamma},$ (17)
[…]
The radius $r$ in Eq. 17 represents the minimum scope within which particles can be covered with a probability close to $1$ during the diffusion process. After choosing the radius $r$, we use the neighbouring grid points to calculate the corresponding mathematical expectation via Eq. 15, which can be accurately approximated with minimal computational cost when $\sum_{i}p_{t}(\boldsymbol{\xi}_{p,t}^{d,i})\delta\approx 1$. However, when the grid size is close to the radius $r$, we may encounter $\sum_{i}p_{t}(\boldsymbol{\xi}_{p,t}^{d,i})\delta\ll 1$ because the random walks of particles are concentrated around $\boldsymbol{\xi}_{p,t}^{c}$ while the grid points are not dense enough. To address this issue, we interpolate the coordinate system $\{\boldsymbol{x}_{p}\}_{p=1}^{P}$ to a high-resolution one $\{\boldsymbol{\bar{x}}_{p}\}_{p=1}^{\bar{P}}$ (yellow points in Fig. 3.B), such that $\sum_{i}p_{t}(\boldsymbol{\bar{\xi}}_{p,t}^{d,i})\bar{\delta}\approx 1$, where $\bar{\delta}$ is the volume of each cell in the high-resolution coordinate system and $\boldsymbol{\bar{\xi}}_{p,t}^{d,i}$ located in the neighbourhood $\bar{N}(\boldsymbol{\xi}_{p,t}^{c},r)$ defined as:
$\bar{N}(\boldsymbol{\xi}_{p,t}^{c},r)\triangleq\{\boldsymbol{\bar{\xi}}_{p,t}^{d,i}:\|\boldsymbol{\bar{\xi}}_{p,t}^{d,i}-\boldsymbol{{\xi}}_{p,t}^{c}\|_{2}\leq r,\boldsymbol{\bar{\xi}}_{p,t}^{d,i}\in\{\boldsymbol{\bar{x}}_{p}\}_{p=1}^{\bar{P}}\}.$ (18)
After that, we approximate the expectation in Eq. 7 in the high-resolution coordinate system as in Eq. 15. Please note that we only have access to the value of $\mathcal{G}_{\theta}(u_{0},t)$ in the low-resolution coordinate system directly, and thus we need to interpolate the physical field $\mathcal{G}_{\theta}(u_{0},t)$ to the corresponding resolution. In practice, we utilize the Fourier transform to map the low-resolution PDE fields to the frequency domain, and use the inverse Fourier transform to remap it to the high-resolution space. We illustrate the choice of $r$ and the interpolation trick of the MCNP Solver in Fig. 3.
4 Experiments
In this section, we conduct numerical experiments to evaluate the proposed MCNP Solver on four tasks: 1D convection-diffusion equation, 1D Allen-Cahn equation, 2D Navier-Stokes equation, and 2D fractional equation on a disk. We introduce the implementation details of each baseline method and the generalization scheme of the test data in Appendix I. We utilize the FNO [43] as the backbone network for MCNP Solver. For each task, we divide the time interval into $10$ uniform frames, and evaluate the model performance via the average relative $\ell_{2}$ error ($E_{\ell_{2}}$) and relative $\ell_{\infty}$ error ($E_{\ell_{\infty}}$) over $10$ frames on 200 test PDE samples. We repeat each experiment with three random seeds in $\{0,1,2\}$ and report the mean value and variance. All experiments are implemented on an NVIDIA RTX 3090 GPU. The source code is publicly available at: https://github.com/optray/MCNP.
4.1 1D Convection-Diffusion Equation
Figure 4: Simulation of 1D convection-diffusion equation. The prediction result (Left) and point-wise error (Right) of MCNP-10 for an example in E1-E6. The x-axis and y-axis represent spatial coordinates and the predicted values (point-wise error).
TABLE I: 1D convection-diffusion equation with varying $\beta$ and $\kappa$. Relative $\ell_{2}$ errors ($E_{\ell_{2}}$), relative $\ell_{\infty}$ errors ($E_{\ell_{\infty}}$) and computational costs for baseline methods and MCNP Solver. The best results are marked in bold, and the second best results are underlined.
Task Model $E_{\ell_{2}}$ (%) $E_{\ell_{\infty}}$ (%) Train Time (H) Param. (M)
E1 $\beta=0.01$ $\kappa=0.001$ PINO-10 0.155 $\pm$0.005 0.157 $\pm$0.010 0.034 0.140
PINO-20 0.075 $\pm$0.001 0.083 $\pm$0.001 0.054 0.140
PINO-100 0.118 $\pm$0.005 0.129 $\pm$0.003 0.257 0.140
PI-DeepONet 0.495 $\pm$0.196 0.705 $\pm$0.312 0.086 0.153
PI-DeepONet-M 0.367 $\pm$0.087 0.594 $\pm$0.092 0.280 0.153
MCNP-10 0.090 $\pm$0.002 0.103 $\pm$0.002 0.032 0.140
MCNP-20 0.119 $\pm$0.010 0.134 $\pm$0.002 0.054 0.140
E2 $\beta=0.01$ $\kappa=0.005$ PINO-10 1.717 $\pm$0.042 1.886 $\pm$0.058 0.034 0.140
PINO-20 0.484 $\pm$0.003 0.536 $\pm$0.002 0.054 0.140
PINO-100 0.179 $\pm$0.027 0.209 $\pm$0.034 0.257 0.140
PI-DeepONet 0.686 $\pm$0.143 0.929 $\pm$0.195 0.086 0.153
PI-DeepONet-M 0.483 $\pm$0.122 0.637 $\pm$0.177 0.280 0.153
MCNP-10 0.128 $\pm$0.014 0.160 $\pm$0.012 0.032 0.140
MCNP-20 0.150 $\pm$0.012 0.180 $\pm$0.014 0.054 0.140
E3 $\beta=0.01$ $\kappa=0.01$ PINO-10 3.387 $\pm$0.031 3.937 $\pm$0.046 0.034 0.140
PINO-20 1.023 $\pm$0.002 1.176 $\pm$0.013 0.054 0.140
PINO-100 0.235 $\pm$0.036 0.273 $\pm$0.036 0.257 0.140
PI-DeepONet 1.268 $\pm$0.140 1.756 $\pm$0.192 0.086 0.153
PI-DeepONet-M 0.659 $\pm$0.069 0.885 $\pm$0.088 0.280 0.153
MCNP-10 0.170 $\pm$0.012 0.210 $\pm$0.009 0.032 0.140
MCNP-20 0.228 $\pm$0.029 0.272 $\pm$0.035 0.054 0.140
E4 $\beta=0.1$ $\kappa=0.001$ PINO-10 6.287 $\pm$0.155 6.783 $\pm$0.354 0.034 0.140
PINO-20 1.567 $\pm$0.120 1.664 $\pm$0.121 0.054 0.140
PINO-100 0.257 $\pm$0.017 0.295 $\pm$0.031 0.257 0.140
PI-DeepONet 0.615 $\pm$0.048 0.889 $\pm$0.078 0.086 0.153
PI-DeepONet-M 0.300 $\pm$0.042 0.376 $\pm$0.046 0.280 0.153
MCNP-10 0.187 $\pm$0.007 0.228 $\pm$0.012 0.032 0.140
MCNP-20 0.201 $\pm$0.017 0.235 $\pm$0.023 0.054 0.140
E5 $\beta=0.1$ $\kappa=0.005$ PINO-10 3.919 $\pm$0.025 4.526 $\pm$0.155 0.034 0.140
PINO-20 1.106 $\pm$0.010 1.318 $\pm$0.023 0.054 0.140
PINO-100 0.222 $\pm$0.015 0.268 $\pm$0.017 0.257 0.140
PI-DeepONet 0.706 $\pm$0.113 0.968 $\pm$0.114 0.086 0.153
PI-DeepONet-M 0.588 $\pm$0.038 0.718 $\pm$0.043 0.280 0.153
MCNP-10 0.172 $\pm$0.018 0.222 $\pm$0.018 0.032 0.140
MCNP-20 0.185 $\pm$0.010 0.235 $\pm$0.013 0.054 0.140
E6 $\beta=0.1$ $\kappa=0.01$ PINO-10 4.784 $\pm$0.031 5.868 $\pm$0.072 0.034 0.140
PINO-20 1.435 $\pm$0.005 1.720 $\pm$0.008 0.054 0.140
PINO-100 0.328 $\pm$0.065 0.385 $\pm$0.063 0.257 0.140
PI-DeepONet 1.440 $\pm$0.410 1.921 $\pm$0.424 0.086 0.153
PI-DeepONet-M 0.827 $\pm$0.051 1.129 $\pm$0.042 0.280 0.153
MCNP-10 0.200 $\pm$0.012 0.259 $\pm$0.012 0.032 0.140
MCNP-20 0.248 $\pm$0.025 0.307 $\pm$0.024 0.054 0.140
In this section, we conduct experiments on periodical 1D convection-diffusion equation defined as follows:
$\frac{\partial u(x,t)}{\partial t}=\beta\nabla u(x,t)+\kappa\Delta u(x,t),\ x\in[0,1],t\in[0,2].$ (19)
The initial states $u(x,0)$ are generated from the functional space $\mathcal{F}_{N}\triangleq\{\sum_{n=1}^{N}a_{n}\sin(2\pi nx):a_{n}\sim\mathbb{U}(0,1)\},$ where $\mathbb{U}(0,1)$ denotes the uniform distribution over $(0,1)$, and $N$ represents the maximum frequency of the functional space.
4.1.1 Experimental Settings
In this setting, $\beta$ and $\kappa$ respectively represent the convection and diffusion rate, which dominate the inherent property of corresponding physical revolution for Eq. 19. To systematically evaluate the performance of the methods employed, we select two different $\beta$ in $\{0.01,0.1\}$ and three different $\kappa$ in $\{0.001,0.005,0.01\}$. These six experimental settings are denoted E1-E6 with $(\beta,\kappa)=(0.01,0.001),(0.01,0.005),(0.01,0.01),(0.1,0.001),\\ (0.1,0.005),(0.1,0.01)$, respectively. In this section, we set the maximum frequency $N$ as 5. We divide the spatial domain $[0,1)$ into $64$ grid elements for all experiments.
4.1.2 Baselines
We introduce the neural PDE solvers performed on 1D convection-diffusion equations, including: i). PINO [45]: an unsupervised neural operator. To evaluate the performance of PINO with varying step sizes, we divide the time interval into 10/20/100 uniform frames, denoted as PINO-10/20/100, respectively. The loss function is constructed through the pseudo-spectral method due to the periodic boundary condition; thus, the boundary conditions can be naturally satisfied. ii). PI-DeepONet [90]: an unsupervised neural operator based on PINN loss and DeepONet, which utilizes the residual of PDE to construct the loss function. iii). PI-DeepONet-M [91]: a modified version of PI-DeepONet, which utilizes an adaptive re-weighting scheme to balance training samples and loss functions. iv). MCNP Solver, we divide the time interval into 10/20 uniform frames, which are denoted as MCNP-10/20, respectively.
4.1.3 Results
Fig. 4 illustrates the predicted $u$ for MCNP-10 from $t=0.2$ to $t=2.0$ for E1-E6, respectively. When comparing the cases with $\beta=0.01$ and $0.1$, the effect of material advection becomes more remarkable with larger $\beta$, with peaks and troughs of the heat flows more noticeably shifting towards the negative $x$-axis direction. Keeping $\beta$ as a constant, as $\kappa$ increases, we observe that the diffusion effect gradually governs the motion of the material, causing the peaks and troughs to decay rapidly. In particular, E4 exhibits the most significant convection effect due to the Péclet number ($\operatorname{Pe}\propto\beta/\kappa$) reaching maximum in this example. When $\operatorname{Pe}$ number is high, the convection effect is dominant, and the diffusion effect can be disregarded, which means that the macroscopic motion of the heat flow has a much greater impact than the random motion of particles. Table I presents each method’s performance and computational cost in the 1D convection-diffusion equation. Among all unsupervised neural PDE solvers, including PI-DeepONet-(M) and PINO-10/20/100, the MCNP-10 performs the best on most tasks, particularly for cases with large convection or diffusion rates. For PINO, its accuracy is not robust with step size, indicating that the pseudo-spectral method cannot provide sufficiently accurate training targets with a coarser partition of the time interval. Additionally, the accuracy of PINO in E4 declines most rapidly with increasing temporal step size. Such phenomenons happen because the $\operatorname{Pe}$ number in this experiment is significantly higher than in other experiments; hence, the conventional Eulerian method requires a finer time step to maintain simulation accuracy. For PINO-100, despite achieving comparable results on E1 and E5, its training cost is eight times more than that of MCNP-10. It is also interesting to note that MCNP-10 performs better than MCNP-20 in Table I. The underlying reason is that the drift coefficient $\beta$ and the forcing $f$ are constants in Eq. 19, and thus, the discrete errors can be eliminated when simulating the trajectories of particles for Feynman-Kac law. As a result, the setting $\Delta t=2/10$ is already sufficient to provide accurate training signals for the MCNP Solver, while further refining the step size would only add additional fitting and generalization burdens.
4.2 1D Allen-Cahn Equation
In this section, we conduct experiments on the 1D Allen-Cahn equation with Dirichlet/Neumann boundary conditions as follows:
$\frac{\partial u(x,t)}{\partial t}=$ $\epsilon\Delta u(x,t)+u(x,t)-u(x,t)^{3},$ (20)
$x\in[0,1],t\in[0,1];$
$\operatorname{Dirichlet:}$ $\ u(0,t)=u(1,t)=0;$
$\operatorname{Neuman:}$ $\ \frac{\partial u(x,t)}{\partial x}\bigg{|}_{x=0}=\frac{\partial u(x,t)}{\partial x}\bigg{|}_{x=1}=0;$
The initial states $u(x,0)$ are generated from the functional space $\mathcal{F}_{N}\triangleq\{\sum_{n=1}^{N}a_{n}\sin(2\pi nx):a_{n}\sim\mathbb{U}(0,1)\}$, and $N$ represents the maximum frequency of the functional space.
4.2.1 Experimental Settings
Firstly, we fix $\epsilon$ as 0.01, and select two different $N$ in $\{5,10\}$ and two kinds of boundary conditions to evaluate the performance of different methods in handling spatial variations and varying boundary conditions. Secondly, when $\epsilon$ tends to zero, the interface layers become thinner and sharper, leading to long-lived metastable structures [15]. To observe such phenomena, we choose $\epsilon$ as 0.0001 with $N$ in $\{5,10\}$, respectively. Therefore, these six experimental settings are denoted E1-E6 with $(\epsilon,N,\operatorname{Boundary})=(0.01,5,\operatorname{D}),(0.01,10,\operatorname{D}),(0.01,5,\operatorname{N}),\\ (0.01,10,\operatorname{N}),(0.0001,5,\operatorname{D}),(0.0001,10,\operatorname{D})$111”D” represents the Dirichlet boundary condition and ”N” represents the Neumann boundary condition., respectively. We divide the spatial domain $[0,1]$ into $65$ uniform grid elements for all experiments.
4.2.2 Baselines
We introduce the neural PDE solvers performed on the 1D Allen-Cahn equations, including: i). PINO [45]: an unsupervised neural operator. We divide the time interval into 20/100 uniform frames, denoted as PINO-20/100. The loss function is constructed through the finite difference method, and an additional loss term is involved in enforcing the neural PDE solver that satisfies the boundary conditions. ii). PI-DeepONet [90]: an unsupervised neural operator based on PINN loss and DeepONets. iii). PI-DeepONet-M [91]: a modified version of PI-DeepONet. iv). MCNP Solver, we divide the time interval into 20/100 uniform frames, denoted as MCNP-20/100, respectively. When applying Feynman-Kac law to Allen-Cahn equation, the term $u-u^{3}$ in Eq. 20 is regarded as the external forcing in Eq. 7.
4.2.3 Results
Figure 5: Simulation of 1D Allen-Cahn equation. The prediction result (Left) and point-wise error (Right) of MCNP-100 for an example in E1-E6. The x-axis and y-axis represent spatial coordinates and the predicted values (point-wise error).
Fig. 5 illustrates the predicted $u$ for MCNP-100 from $t=0.1$ to $t=1.0$ for E1 through E6, respectively. The simulation of the Allen-Cahn equation is more challenging than the convection-diffusion equation due to its nonlinearity. We can observe that higher $N$ values can result in more obvious spatial variations, particularly near the initial time. In the experiments with Dirichlet boundary conditions, the system values decay rapidly near the boundary, whereas in the case of Neumann conditions, they maintain a relatively stable variation. From a microscopic perspective, this occurs because the particles are absorbed after hitting the Dirichlet boundary, while they are reflected after colliding with the Neumann boundary. Unlike other baseline methods, MCNP Solver does not rely on additional loss functions to encode the boundary conditions. Instead, the MCNP Solver can naturally satisfy these boundary conditions by choosing the corresponding transition probability during its random walk. The simulation results in Fig. 5 demonstrate that the MCNP Solver effectively meets the Dirichlet and Neumann boundary conditions when solving the Allen-Cahn equations. It can be seen in Fig. 5 that when $\epsilon=0.0001$, the phase field $u(x)$ oscillates rapidly between positive and negative values, highlighting the sharpness of the interface layers due to the very small $\epsilon$. These oscillations are caused by the equation striving to minimize the energy, leading to high-frequency changes in the solution near the interface to accommodate the small $\epsilon$ [15]. Table II presents each method’s performance and computational cost on the 1D Allen-Cahn equation. The results of PI-DeepONet indicate that the PINN loss cannot efficiently handle high-frequency components, which has also been observed in previous literature [36, 92]. Although PI-DeepONet-M mitigates this issue via adaptive sampling, there is still a gap in precision when solving high-frequency problems compared to other methods. Among all unsupervised methods, only MCNP-100 achieves a relative error lower than 1% on all experiments and metrics. Moreover, the performance of MCNP-20 is generally comparable to that of PINO-100 while taking only around 21% of the training time, demonstrating the advantages of neural Monte Carlo loss to learn PDEs unsupervised.
TABLE II: 1D Allen-Cahn equation with varying $\epsilon$, $N$ and boundary conditions. Relative $\ell_{2}$ errors ($E_{\ell_{2}}$), relative $\ell_{\infty}$ errors ($E_{\ell_{\infty}}$) and computational costs for baseline methods and MCNP Solver. The best results are marked in bold, and the second best results are underlined.
Task Model $E_{\ell_{2}}$ (%) $E_{\ell_{\infty}}$ (%) Train Time (H) Param. (M)
E1 $\epsilon=0.01$ $N=5$ Dirichlet PINO-20 1.707 $\pm$0.262 2.082 $\pm$0.271 0.135 0.140
PINO-100 0.639 $\pm$0.020 0.736 $\pm$0.009 0.627 0.140
PI-DeepONet 2.444 $\pm$1.022 3.551 $\pm$1.579 0.431 0.219
PI-DeepONet-M 1.071$\pm$ 0.113 1.431 $\pm$0.125 1.004 0.219
MCNP-20 0.918 $\pm$0.037 1.108 $\pm$0.034 0.133 0.140
MCNP-100 0.547 $\pm$0.053 0.636 $\pm$0.061 0.621 0.140
E2 $\epsilon=0.01$ $N=10$ Dirichlet PINO-20 3.180 $\pm$0.182 4.534 $\pm$0.232 0.135 0.140
PINO-100 1.165 $\pm$0.192 1.430 $\pm$0.193 0.627 0.140
PI-DeepONet 5.615 $\pm$1.191 7.163 $\pm$1.455 0.431 0.219
PI-DeepONet-M 2.659 $\pm$0.259 3.241 $\pm$0.240 1.004 0.219
MCNP-20 1.290 $\pm$0.062 1.592 $\pm$0.079 0.133 0.140
MCNP-100 0.831 $\pm$0.200 0.955 $\pm$0.212 0.621 0.140
E3 $\epsilon=0.01$ $N=5$ Neumann PINO-20 1.727 $\pm$0.368 2.885 $\pm$0.043 0.136 0.140
PINO-100 1.705 $\pm$0.019 2.768 $\pm$0.043 0.627 0.140
PI-DeepONet 1.773 $\pm$0.890 2.353 $\pm$1.128 0.433 0.219
PI-DeepONet-M 0.438 $\pm$0.047 0.617 $\pm$0.062 1.012 0.219
MCNP-20 0.598 $\pm$0.011 0.801 $\pm$0.018 0.133 0.140
MCNP-100 0.394 $\pm$0.037 0.518 $\pm$0.040 0.622 0.140
E4 $\epsilon=0.01$ $N=10$ Neumann PINO-20 2.178 $\pm$0.368 2.885 $\pm$0.368 0.136 0.140
PINO-100 1.176 $\pm$0.025 1.805 $\pm$0.043 0.627 0.140
PI-DeepONet 2.587 $\pm$0.247 3.111 $\pm$0.300 0.433 0.219
PI-DeepONet-M 2.169 $\pm$0.072 2.489 $\pm$0.091 1.012 0.219
MCNP-20 1.218 $\pm$0.016 1.470 $\pm$0.026 0.133 0.140
MCNP-100 0.582 $\pm$0.078 0.671 $\pm$0.083 0.622 0.140
E5 $\epsilon=0.0001$ $N=5$ Dirichlet PINO-20 0.735 $\pm$0.162 1.445 $\pm$0.275 0.135 0.140
PINO-100 0.330 $\pm$0.004 0.842 $\pm$0.003 0.627 0.140
PI-DeepONet 8.251 $\pm$2.665 15.853 $\pm$4.519 0.431 0.219
PI-DeepONet-M 3.781$\pm$ 1.943 4.724 $\pm$2.568 1.004 0.219
MCNP-20 0.285 $\pm$0.010 0.593 $\pm$0.014 0.156 0.140
MCNP-100 0.303 $\pm$0.018 0.594 $\pm$0.047 0.651 0.140
E6 $\epsilon=0.0001$ $N=10$ Dirichlet PINO-20 3.605 $\pm$1.030 7.181 $\pm$1.636 0.135 0.140
PINO-100 1.817 $\pm$0.008 4.114 $\pm$0.042 0.627 0.140
PI-DeepONet 15.009 $\pm$2.398 25.870 $\pm$3.776 0.431 0.219
PI-DeepONet-M 9.542$\pm$ 1.058 12.357 $\pm$1.569 1.004 0.219
MCNP-20 1.459 $\pm$0.007 2.957 $\pm$0.011 0.156 0.140
MCNP-100 1.487 $\pm$0.090 3.083 $\pm$0.230 0.651 0.140
4.3 2D Navier-Stokes Equation
In this experiment, we simulate the vorticity field for 2D incompressible flows in a periodic domain $\Omega=[0,1]\times[0,1]$, whose vortex equation is given as follows:
$\frac{\partial\omega}{\partial t}$ $=-(\boldsymbol{u}\cdot\nabla)\omega+\nu\Delta\omega+f(\boldsymbol{x}),\ \boldsymbol{x}\in\Omega,t\in[0,10],$ (21)
$\omega$ =×$=\nabla\times\boldsymbol{u},$
where $f(\boldsymbol{x})$ is the external forcing, $\boldsymbol{u}\in\mathbb{R}^{2}$ denotes the velocity, and $\nu\in\mathbb{R}^{+}$ represents the viscosity coefficient. The initial vorticity is generated from the Gaussian random field $\mathcal{N}\left(0,7^{3/2}(-\Delta+49\boldsymbol{I})^{-2.5}\right)$ with periodic boundaries. In this section, we consider three kinds of external forcing, including: (1) zero forcing $f_{1}(\boldsymbol{x})\equiv 0$; (2) Li forcing $f_{2}(\boldsymbol{x})\triangleq 0.1\sin\left(2\pi\left(\boldsymbol{x}_{1}+\boldsymbol{x}_{2}\right)\right)+0.1\cos\left(2\pi\left(\boldsymbol{x}_{1}+\boldsymbol{x}_{2}\right)\right)$, which is a classical forcing in the paper of FNO [43]; (3) Kolmogorov forcing $f_{3}(\boldsymbol{x})\triangleq 0.1\cos(8\pi\boldsymbol{x}_{1})$, which can result in a much wider range of trajectories due to the chaotic [81].
4.3.1 Experimental Settings
The viscosity coefficient $\nu$ can be regarded as a measure of the spatiotemporal complexity of the Navier-Stokes equation. As $\nu$ decreases, the nonlinear term $(\boldsymbol{u}\cdot\nabla)\omega$ gradually governs the motion of fluids, increasing the difficulty of simulation. To evaluate the performance of handling different degrees of turbulence, we conduct the experiments with $\nu$ in $\{10^{-4},10^{-5}\}$. Therefore, six experimental settings are denoted E1-E6 with $(f(\boldsymbol{x}),\nu)=(f_{1}(\boldsymbol{x}),10^{-4}),(f_{1}(\boldsymbol{x}),10^{-5}),(f_{2}(\boldsymbol{x}),10^{-4}),(f_{2}(\boldsymbol{x}),10^{-5}),\\ (f_{3}(\boldsymbol{x}),10^{-4}),(f_{3}(\boldsymbol{x}),10^{-5})$, respectively. We divide the domain $\Omega$ into $64\times 64$ uniform grid elements.
Figure 6: Simulation of 2D Navier-Stokes equation. The prediction result (Left) and point-wise error (Right) of MCNP-20 for an example in E1-E6.
TABLE III: 2D Navier-Stokes equation with varying $\nu$ and forcing. Relative $\ell_{2}$ errors ($E_{\ell_{2}}$), relative $\ell_{\infty}$ errors ($E_{\ell_{\infty}}$) and computational costs for baseline methods and MCNP Solver. The best results are marked in bold, and the second best results are underlined. The training time for MCNP varies among the six experiments due to the interpolation trick.
Task Model $E_{\ell_{2}}$ (%) $E_{\ell_{\infty}}$ (%) Train Time (H) Param. (M)
E1 $\nu=10^{-4}$ Zero PINO-10 3.693 $\pm$0.081 6.195 $\pm$0.084 0.232 5.319
PINO-20 2.749 $\pm$0.091 4.654 $\pm$0.129 0.413 5.319
PINO-100 3.027 $\pm$0.134 5.155 $\pm$0.188 1.919 5.319
SP-FNO-10 4.348 $\pm$0.077 7.597 $\pm$0.084 $>12$ 5.319
SP-FNO-20 3.821 $\pm$0.264 6.855 $\pm$0.337 $>24$ 5.319
MCNP-10 2.397 $\pm$0.101 4.102 $\pm$0.139 0.256 5.319
MCNP-20 2.680 $\pm$0.130 4.565 $\pm$0.202 0.488 5.319
E2 $\nu=10^{-5}$ Zero PINO-10 10.035 $\pm$0.229 14.685 $\pm$0.250 0.232 5.319
PINO-20 8.464 $\pm$0.402 12.596 $\pm$0.441 0.413 5.319
PINO-100 7.554 $\pm$0.206 11.627 $\pm$0.384 1.919 5.319
SP-FNO-10 9.023 $\pm$0.154 14.487 $\pm$0.208 $>12$ 5.319
SP-FNO-20 8.677 $\pm$0.090 13. 217$\pm$0.183 $>24$ 5.319
MCNP-10 7.055 $\pm$0.123 10.815 $\pm$0.138 0.256 5.319
MCNP-20 7.620 $\pm$0.184 11.564 $\pm$0.201 0.527 5.319
E3 $\nu=10^{-4}$ Li PINO-10 5.502 $\pm$0.040 9.002 $\pm$0.056 0.232 5.319
PINO-20 3.971 $\pm$0.115 6.819 $\pm$0.184 0.413 5.319
PINO-100 3.366 $\pm$0.034 5.908 $\pm$0.041 1.919 5.319
SP-FNO-10 3.794$\pm$ 0.076 6.885 $\pm$0.109 $>12$ 5.319
SP-FNO-20 3.638 $\pm$0.056 6.516 $\pm$0.082 $>24$ 5.319
MCNP-10 3.101 $\pm$0.043 5.815 $\pm$0.075 0.261 5.319
MCNP-20 2.999 $\pm$0.082 5.336 $\pm$0.131 0.491 5.319
E4 $\nu=10^{-5}$ Li PINO-10 10.617 $\pm$0.166 16.799 $\pm$0.255 0.232 5.319
PINO-20 8.124 $\pm$0.261 13.636 $\pm$0.351 0.413 5.319
PINO-100 6.439 $\pm$0.105 11.123 $\pm$0.139 1.919 5.319
SP-FNO-10 6.561 $\pm$0.098 11.529$\pm$ 0.134 $>12$ 5.319
SP-FNO-20 6.320 $\pm$0.109 11.214 $\pm$0.144 $>24$ 5.319
MCNP-10 5.825 $\pm$0.068 10.452 $\pm$0.122 0.260 5.319
MCNP-20 5.776 $\pm$0.085 10.137 $\pm$0.146 0.531 5.319
E5 $\nu=10^{-4}$ Kolmogorov PINO-10 6.652 $\pm$0.141 8.922 $\pm$0.205 0.232 5.319
PINO-20 5.164 $\pm$0.138 7.195 $\pm$0.196 0.413 5.319
PINO-100 4.925 $\pm$0.037 6.972 $\pm$0.042 1.919 5.319
SP-FNO-10 6.801 $\pm$0.124 9.581 $\pm$0.272 $>12$ 5.319
SP-FNO-20 6.682 $\pm$0.040 9.525 $\pm$0.049 $>24$ 5.319
MCNP-10 4.799 $\pm$0.114 6.443 $\pm$0.086 0.278 5.319
MCNP-20 4.609 $\pm$0.048 6.488 $\pm$0.053 0.518 5.319
E6 $\nu=10^{-5}$ Kolmogorov PINO-10 16.342 $\pm$0.345 22.497 $\pm$0.266 0.232 5.319
PINO-20 13.501 $\pm$0.446 19.200 $\pm$0.748 0.413 5.319
PINO-100 11.874 $\pm$0.235 16.973 $\pm$0.204 1.919 5.319
SP-FNO-10 13.789 $\pm$0.429 19.004 $\pm$0.517 $>12$ 5.319
SP-FNO-20 13.601 $\pm$0.390 19.214 $\pm$0.445 $>24$ 5.319
MCNP-10 10.161 $\pm$0.150 15.053 $\pm$0.227 0.252 5.319
MCNP-20 10.829 $\pm$0.154 15.764 $\pm$0.196 0.528 5.319
4.3.2 Baselines
We introduce the baselines conducted on 2D Navier-Stokes equation, including: [^2] For PI-DeepONets [90, 91], they only conduct experiments on time-independent PDE in 2D situations in their paper. i). PINO [45]: we divide the time interval $[0,10]$ into 10/20/100 uniform frames, denoted as PINO-10/20/100, respectively. The loss function is constructed through the pseudo-spectral method due to the periodic boundary conditions. ii). SP-FNO  [56, 57], an unsupervised operator learning method that employs stochastic projection (SP) to enhance the precision of spatial gradient computation. We use FNO as the backbone model for a fair comparison and divide the time interval into uniform 10/20 frames, denoted as SP-FNO-10/20. iii). MCNP Solver, we divide the time interval into uniform 10/20 frames, denoted as MCNP-10/20, respectively. When applying Feynman-Kac law to Navier-Stokes equation, the velocity $\boldsymbol{u}$ in Eq. 21 is regarded as the drift term $\boldsymbol{\beta}$ in Eq. 7.
4.3.3 Results
Fig. 6 shows the predicted vorticity field $\omega$ of a learned MCNP Solver from $t=2$ to $t=10$ in E1-E6, respectively. Compared to the case with $\nu=10^{-4}$, we observe that the vorticity field has more remarkable peak and trough values with more intricate details when $\nu=10^{-5}$, which implies that the fluid elements are rotating at higher speeds. This instability is caused by the nonlinear convection terms gradually taking control of the motion of fluids. Additionally, the external forcing $f$ significantly influences the trend of fluid motion. When the external forcing is zero, the fluid motion primarily arises from internal dynamical processes within the system, such as the interaction between viscous and inertial forces, resulting in a relatively slow change in the vorticity field over time. When an external forcing drives the system, the corresponding vorticity field changes markedly in response to the forcing, leading to a faster evolution of the vorticity field. In particular, when driven by a Kolmogorov forcing, numerous new vortices of different scales are generated during the flow, making the fluid simulation more challenging. More simulation results of other baseline methods can be seen in the Appendix  V. Table III presents each method’s performance and computational cost on the 2D Navier-Stokes equations. Compared to other unsupervised training methods, the MCNP Solver achieved the lowest relative error for all tasks and metrics. Unlike PINO, whose performance is highly affected by step size, especially for cases with $\nu=10^{-5}$, the MCNP Solver can maintain stability with respect to step size. It is worth mentioning that MCNP-10 outperforms PINO-100 on all metrics while only taking 13.1-14.5% of the training time. Such an advantage stems from the fact that the Lagrangian method can be better adapted to coarse step size, as discussed in [29, 35].
4.4 2D Fractional Diffusion Equation on a Disk
In this section, we simulate the fractional diffusion equation on a 2D unit circle as follows:
${\frac{\partial u}{\partial t}}$ ${=-\kappa(-\Delta)^{\alpha/2}u(\boldsymbol{x},t),\quad\|\boldsymbol{x}\|_{2}\leq 1,t\in[0,1],}$ (22)
${u(\boldsymbol{x},t)}{=0\quad\text{for}\ \|\boldsymbol{x}\|_{2}=1,}$
where $(-\Delta)^{\alpha/2}$ denotes the fractional Laplacian operator, defined via the following hyper-singular integral [46]:
${(-\Delta)^{\alpha/2}u(\boldsymbol{x})\triangleq C_{\alpha}\text{ P.V. }\int_{\mathbb{R}^{2}}\frac{u(\boldsymbol{x})-u(\boldsymbol{y})}{\|\boldsymbol{x}-\boldsymbol{y}\|_{2}^{2+\alpha}}\mathrm{d}\boldsymbol{y},\ 0<\alpha<2,}$ (23)
where P.V. denotes the principle value and $C_{\alpha}$ is defined as:
${C_{\alpha}=\frac{2^{\alpha}\Gamma\left(\frac{\alpha+2}{2}\right)}{\pi|\Gamma(-\alpha/2)|}.}$ (24)
As shown in Eq. 22, calculating the fractional Laplacian operator is challenging due to its singularity and non-local properties. However, we can efficiently simulate it from the probability perspective, where we only need to replace the Brownian motion in Eq. 4 with the $\alpha$-stable Lévy process [34, 99, 100]. The initial states $u(\boldsymbol{x},0)$ are generated from the functional space $\mathcal{H}_{N}\triangleq\{\sum_{n=1}^{N}a_{n}(1-\|\boldsymbol{x}\|_{2}^{2})^{n+1}:a_{n}\sim\mathbb{U}(0,1)\}$, and $N$ represents the maximum degree of the functional space.
4.4.1 Experimental Settings
We select four different $\alpha$ in $\{0.5,1.0,1.5,2.0\}$ to evaluate the performance of MCNP Solver in handling varying fractional coefficients. These four experiments are denoted as E1-E4. We divide the time interval $[0,1]$ into $100$ uniform intervals. We set the maximum degree $N$ as 10. Because Eq. 22 is defined on a disk, we use DeepONets as backbone models and train MCNP Solver in a mesh-free regime.
Figure 7: Simulation of 2D fractional diffusion equation on a disk. The prediction result (Above) and point-wise error (Below) of MCNP Solver for an example in E1-E4 at $t=1.0$.
TABLE IV: 2D fractional diffusion equation with varying $\alpha$. Relative $\ell_{2}$ errors ($E_{\ell_{2}}$), relative $\ell_{\infty}$ errors ($E_{\ell_{\infty}}$) and computational costs.
Task $E_{\ell_{2}}$ (%) $E_{\ell_{\infty}}$ (%) Train Time (H) Param. (M)
E1 ($\alpha=0.5$) 1.139 $\pm$0.145 1.689 $\pm$0.320 0.139 0.325
E2 ($\alpha=1.0$) 1.538 $\pm$0.616 1.846 $\pm$1.015 0.139 0.325
E3 ($\alpha=1.5$) 1.424 $\pm$0.266 1.386 $\pm$0.315 0.139 0.325
E4 ($\alpha=2.0$) 0.977 $\pm$0.045 1.300 $\pm$0.132 0.139 0.325
4.4.2 Results
Fig. 7 shows the predicted physical field $u$ of a learned MCNP Solver at $t=1.0$ in E1-E4, respectively. Comparing different $\alpha$ values, we observe that the diffusion effect becomes more pronounced as the fractional coefficient $\alpha$ increases. Table IV presents each method’s performance and computational cost on the 2D fractional diffusion equation. The results reveal that the MCNP Solver has consistently maintained an error rate under 2% for all tasks and performance indicators. It is interesting to note that the precision of the simulation does not vary monotonically with changes in the fractional coefficient $\alpha$. This phenomenon happens because excessively large or small values of $\alpha$ present challenges to the simulation tasks. For smaller $\alpha$ values, the singularities of the fractional operators would become more remarkable. From a microscopic perspective, particles are more likely to undergo significant jumps following a Lévy process. Conversely, the diffusion rate accelerates with larger $\alpha$ values, leading to a more rapid evolution of the entire physical field. It is worth mentioning that despite the non-rectangular geometry in this section, the MCNP Solver can be extended to a mesh-free regime and preserve the boundary conditions automatically through the random walk of particles.
5 Additional Experimental Results
In this section, we conduct several additional numerical experiments to comprehensively compare the scopes and advantages of the proposed MCNP Solver with other widely-used PDE solvers, including the Monte Carlo methods, traditional Eulerian methods and supervised solver learning methods. Furthermore, we conduct ablation studies on MCNP Solver to study the effects of the proposed tricks.
TABLE V: In comparing the relative $\ell_{2}$ errors (%) between MCNP and other numerical solvers across four types of equations. In column 9, we report the inference time (S), denoted as $T_{\operatorname{GPU}}/T_{\operatorname{CPU}}$, for a single test sample on both GPU and CPU. The inference time for the fractional PDE on the GPU is not included, as its implementation is based on a CPU package. In column 10, we report the training time of MCNP for each experiment.
Equation Solver E1 E2 E3 E4 E5 E6 $T_{\operatorname{GPU}}/T_{\operatorname{CPU}}$ (S) Train Time (H)
Convection- Diffusion MCM 3.374 $\pm$0.047 4.648 $\pm$0.129 7.269 $\pm$0.382 7.269 $\pm$0.382 4.612 $\pm$0.098 7.248 $\pm$0.301 0.163 / 0.236
MCM+ 1.932 $\pm$0.008 1.440 $\pm$0.050 2.144 $\pm$0.045 6.952 $\pm$0.010 1.427 $\pm$0.031 2.097 $\pm$0.065 0.188 / 0.518
MCM++ 1.718 $\pm$0.003 0.476 $\pm$0.013 0.718 $\pm$0.036 6.896 $\pm$0.010 0.489 $\pm$0.011 0.737 $\pm$0.022 0.279 / 3.045
PSM 0.115 1.121 2.640 3.278 2.428 3.686 0.005 / 0.003
PSM+ 0.056 0.312 0.644 0.733 0.588 0.843 0.007 / 0.005
PSM++ 0.043 0.101 0.165 0.184 0.160 0.209 0.014 / 0.010
MCNP 0.090 $\pm$0.002 0.128 $\pm$0.014 0.170 $\pm$0.012 0.187 $\pm$0.007 0.172 $\pm$0.018 0.200 $\pm$0.012 0.002 / 0.009 0.032-0.054
Allen- Cahn MCM 7.430 $\pm$0.102 7.458 $\pm$0.096 16.155 $\pm$0.343 18.133 $\pm$0.459 7.368 $\pm$0.203 12.068 $\pm$0.023 0.048 / 0.257
MCM+ 2.902 $\pm$0.059 2.992 $\pm$0.071 8.772 $\pm$0.570 10.079 $\pm$0.521 3.449 $\pm$0.048 10.821 $\pm$0.068 0.063 / 2.692
MCM++ 2.103 $\pm$0.014 2.129 $\pm$0.028 1.265 $\pm$0.002 1.848 $\pm$0.005 2.590 $\pm$0.022 7.110 $\pm$0.040 0.551 / 14.568
FDM NAN NAN NAN NAN 1.812 12.441 0.005 / 0.002
FDM+ 0.983 1.468 1.256 2.166 0.480 1.987 0.021 / 0.009
FDM++ 0.625 0.788 1.424 1.140 0.366 1.886 0.043 / 0.018
MCNP 0.547 $\pm$0.053 0.831 $\pm$0.200 0.394 $\pm$0.037 0.582 $\pm$0.078 0.285 $\pm$0.010 1.487 $\pm$0.090 0.002 / 0.009 0.133-0.651
Navier- Stokes MCM 7.326 $\pm$0.035 7.452 $\pm$0.045 5.864 $\pm$0.032 4.884 $\pm$0.012 6.890 $\pm$0.026 6.994 $\pm$0.011 0.180 / 4.552
MCM+ 5.396 $\pm$0.043 7.055 $\pm$0.010 4.523 $\pm$0.026 4.648 $\pm$0.007 5.316 $\pm$0.035 6.700 $\pm$0.014 0.225 / 7.968
MCM++ 3.744 $\pm$0.006 6.814 $\pm$0.005 3.453 $\pm$0.008 4.491 $\pm$0.003 4.023 $\pm$0.025 6.519 $\pm$0.011 0.469 / 16.647
PSM 4.583 12.336 6.116 11.518 4.500 9.920 0.069 / 0.107
PSM+ 4.227 11.427 3.508 8.443 4.122 9.335 0.139 / 0.204
PSM++ 4.102 11.124 2.946 6.789 4.007 9.147 0.344 / 0.513
MCNP 2.397 $\pm$0.101 7.055 $\pm$0.123 2.999 $\pm$0.082 5.776 $\pm$0.085 4.609 $\pm$0.048 10.161 $\pm$0.150 0.007 / 0.065 0.252-0.531
Fractional MCM 5.668 $\pm$0.629 5.710 $\pm$0.886 5.483 $\pm$0.866 6.511 $\pm$0.198 – / 0.058
MCM+ 5.139 $\pm$0.219 5.330 $\pm$0.236 4.074 $\pm$0.106 2.821 $\pm$0.168 – / 0.518
MCM++ 5.057 $\pm$0.066 5.129 $\pm$0.056 3.799 $\pm$0.088 2.015 $\pm$0.010 – / 6.625
PSM 0.114 0.499 1.908 4.644 – / 0.011
PSM+ 0.061 0.252 0.928 2.076 – / 0.018
PSM++ 0.032 0.114 0.378 0.820 – / 0.039
MCNP 1.139 $\pm$0.145 1.538 $\pm$0.616 1.424 $\pm$0.266 0.977 $\pm$0.045 – / 0.010 0.139
5.1 Comparison with Numerical Methods
This section presents experiments comparing the MCNP with various numerical solvers, including the traditional Monte Carlo method and Eulerian methods, such as PSM and FDM, for all equations in Section 4. For Monte Carlo methods, we directly simulate the corresponding SDEs for convection-diffusion and fractional equations, using the branching diffusion method [24] for Allen-Cahn and the random vortex method [66] for Navier-Stokes. To evaluate the effect of particle count, we vary the number of particles, with more plus signs (MCM/MCM+/MCM++) indicating a higher number of particles. For Eulerian solvers, we employ FDM for Allen-Cahn equations and PSM for the other three types of PDEs. In this case, we vary the step size, with more plus signs (FDM/FDM+/FDM++ or PSM/PSM+/PSM++) representing a smaller step size. To ensure a fair comparison, we test the solvers across multiple spatial resolutions, selecting the lowest resolution that yields comparable results. The implementation details are provided in Appendix I.5. Table V summarizes the comparison between MCNP and other numerical solvers.
According to the results in Table V, the performance and efficiency of Monte Carlo methods are usually heavily constrained by the number of particles, particularly at high diffusion rates. For instance, when $\kappa$ is large (as in E3 and E6), there is a significant error difference between MCM and MCM++ in the convection-diffusion equation. While increasing the number of particles can reduce errors, it introduces serious computational challenges, including memory and inference time. In contrast, the MCNP solver approximates the expectation in Eq. 7 using the PDF of neighboring grid points, avoiding the need to sample a large number of particles, as discussed in Section 3.3.2.
In comparing MCNP with traditional Eulerian methods, we observe that conventional solvers often require refining the time step, especially in highly nonlinear cases, leading to increased computational costs. In contrast, MCNP offers significant speed advantages, particularly in solving equations like Navier-Stokes equations, achieving 10-50x speedup on GPU and 2-8x on CPU. Nevertheless, the speedup is limited in some 1D experiments, such as E1 for convection-diffusion, with MCNP sometimes underperforming on the CPU. This suggests that traditional methods can still be more efficient in some simple cases, where the computational complexity is reduced. Another interesting observation is that the GPU provides a more significant speedup for MCNP than for traditional solvers. Such greater GPU speedup is due to MCNP’s ability to leverage spatiotemporal parallelism and output physical fields directly at any time step without sequential iteration. On the other hand, traditional methods often struggle to take full advantage of GPU acceleration, as they are typically designed with sequential computations in the temporal dimension.
Furthermore, like other neural operators, the MCNP Solver requires hyperparameter tuning and training, while traditional numerical methods do not. However, once training is complete, the MCNP Solver can perform inference on new initial value problems without additional training or tuning. This feature makes neural operators particularly promising for applications in inverse design and real-time physical simulations, such as weather forecasting and fluid control [83, 37]. In this section, we report the inference time for the MCNP Solver, while the hyperparameter tuning costs are discussed in Appendix IV.
TABLE VI: Compared to the FNO on 2D Navier-Stokes equation with varying $\nu$. Relative $\ell_{2}$ errors ($E_{\ell_{2}}$), relative $\ell_{\infty}$ errors ($E_{\ell_{\infty}}$) and the training costs for baseline methods and MCNP Solver. The training time of MCNP Solver reported in this table is the average over E3 and E4.
E3 ($\nu=10^{-4}$) E4 ($\nu=10^{-5}$) Time (H)
Method $E_{\ell_{2}}$ (%) $E_{\ell_{\infty}}$ (%) $E_{\ell_{2}}$ (%) $E_{\ell_{\infty}}$ (%) Data Train Total
FNO 4.754 $\pm$0.139 8.934 $\pm$0.248 8.003 $\pm$0.161 15.184 $\pm$0.239 0.346 0.207 0.553
FNO+ 3.460 $\pm$0.163 6.630 $\pm$0.358 6.115 $\pm$0.085 11.743 $\pm$0.143 0.692 0.207 0.899
FNO++ 2.619 $\pm$0.124 4.906 $\pm$0.195 4.640 $\pm$0.032 8.839 $\pm$0.148 1.384 0.207 1.591
MCNP-10 3.101 $\pm$0.043 5.815 $\pm$0.075 5.825 $\pm$0.068 10.452 $\pm$0.122 0 0.261 0.261
MCNP-20 2.999 $\pm$0.082 5.336 $\pm$0.131 5.776 $\pm$0.085 10.137 $\pm$0.146 0 0.511 0.511
5.2 Comparison with Supervised Solver Learning
In this section, we conduct experiments to compare MCNP Solver with supervised neural operator learning methods FNO [43] on the Navier-Stokes equations (E3 and E4) in Section 4.3. To evaluate the performance of FNO with the amounts of datasets, we utilize 500/1000/2000 PDE trajectories to train FNO and denote the corresponding methods as FNO/FNO+/FNO++, respectively.
Table VI presents a comparison between MCNP-10/20 and three versions of FNO. Compared to the supervised methods, which use pre-simulated fixed data for training, MCNP Solver can sample new initial fields per epoch, thereby increasing the diversity of training data. As a result, MCNP-10/20 can outperform FNO and FNO+. With the increase of training data, FNO++ achieves better results. This is because the training data is generated from high-precision numerical methods, while the unsupervised methods construct the loss function with low spatiotemporal resolution. However, as a trade-off, FNO++ spends 211-510% more total time than MCNP-20/10.
TABLE VII: Ablation Studies of each component in MCNP Solver. Relative error (%) and training time for each method on the Navier-Stokes equation tasks with $\nu=10^{-4}$ (E3) and $\nu=10^{-5}$ (E4). The training time of each method reported in this table is the average over E3 and E4.
E3 ($\nu=10^{-4}$) E4 ($\nu=10^{-5}$)
Method $E_{\ell_{2}}$ (%) $E_{\ell_{\infty}}$ (%) $E_{\ell_{2}}$ (%) $E_{\ell_{\infty}}$ (%) Time (H)
MCNP-H-10 11.572 $\pm$0.009 19.482 $\pm$0.013 13.374 $\pm$0.055 23.280 $\pm$0.076 0.253
MCNP-I-10 3.168 $\pm$0.026 5.931 $\pm$0.048 61.719 $\pm$3.475 78.465 $\pm$9.320 0.242
MCNP-10 3.101 $\pm$0.043 5.815 $\pm$0.075 5.825 $\pm$0.068 10.452 $\pm$0.122 0.261
MCNP-H-20 7.104 $\pm$0.048 12.403 $\pm$0.080 9.249 $\pm$0.041 16.613 $\pm$0.101 0.509
MCNP-I-20 8.970 $\pm$0.025 15.138 $\pm$0.026 69.067 $\pm$5.975 72.800 $\pm$4.340 0.408
MCNP-20 2.999 $\pm$0.082 5.336 $\pm$0.131 5.776 $\pm$0.085 10.137 $\pm$0.146 0.511
5.3 Ablation Studies
In this section, we conduct several ablation studies on the MCNP Solver applied to the Navier-Stokes equation (E3 and E4). Our goal is to evaluate the individual contribution of each method component. MCNP-H replaces Heun’s method (Section 3.3.1) with the traditional Euler method when simulating the SDEs. MCNP-I represents the MCNP Solver without the interpolation trick introduced in Section 3.3.3.
Table VII reports the results and training costs. Compared to MCNP Solver with MCNP-H, the results show that using Heun’s method to simulate the SDEs significantly improves the accuracy of MCNP Solver, while incurring only minimal additional computational cost, especially for MCNP-10. Compared to MCNP Solver with MCNP-I, the interpolation trick plays a crucial role in E4 when $\nu=10^{-5}$. As discussed in Section 3.3.3, extremely low diffusion rates can lead to very short-distance diffusion effects, resulting in $\sum_{i}p_{t}(\boldsymbol{\xi}_{p,t}^{d,i})\delta\ll 1$. Therefore, interpolating the original vorticity field becomes essential to ensure that the PDF of grid points satisfies normalization conditions in Eq. 17. Similarly, the interpolation trick can reduce more relative error for MCNP-20 than MCNP-10 in E3 because fine step size can also introduce localized random walks per step.
6 Conclusion and Discussion
In this paper, we propose the MCNP Solver, which leverages the Feynman-Kac formula to train neural PDE solvers in an unsupervised manner. Compared to other unsupervised neural PDE solvers, the MCNP Solver can be more robust to complex spatiotemporal variations due to the advantages of Lagrangian methods. Moreover, the experiments on 2D fractional diffusion equations demonstrate the applicability of the MCNP Solver in mesh-free scenarios and its capability to handle fractional order Laplacian operators.
This paper has several limitations: (1) Some PDEs are not suitable for the Feynman-Kac formula and therefore do not fall within the scope of the MCNP Solver, such as third or higher-order PDEs (involving high-order operators like $u_{xxx}$). (2) The accuracy of the MCNP Solver cannot outperform traditional numerical solvers when disregarding inference time, which is also a major drawback for other existing neural solvers [83, 19], as we discussed in Section 3.3.2. (3) Like other neural operators, the MCNP Solver requires hyperparameter tuning and training, while traditional numerical methods do not. However, once trained, neural operators can be applied directly to new initial value problems for rapid inference.
Furthermore, we suggest several directions for future research: (1) Extend the proposed MCNP Solver to broader scenarios, such as high-dimensional PDEs and optimal control problems; (2) Utilize techniques from out-of-distribution generalization [79] to improve the generalization ability of MCNP Solver; (3) There are some mathematical works have extended the probabilistic representation of PDEs to the higher-order cases [1, 60], and extending the MCNP Solver to such scenario is also a feasible and promising direction.
Acknowledgments
This work was in part supported by NSFC [No. 9247010235], National Key Research and Development Program of China [2019YFA0709501] and CAS Project for Young Scientists in Basic Research [No. YSBR-034].
Appendix  I: Implementation Details
I.1: Baselines
In this paper, we adopt Pytorch [64] to implement MCNP Solver, FNO, SP-FNO, and PINO, and JAX [6] for PI-DeepONet-(M), respectively. Here, we introduce PINO and PI-DeepONet as follows.
PI-DeepONet [90]
PI-DeepONet utilizes the PDE residuals to train DeepONets in an unsupervised way. The loss function in PI-DeepONet can be formulated as follows:
$\mathcal{L}_{\operatorname{PI-DeepONet}}$ $=\lambda_{1}\mathcal{L}_{\operatorname{initial}}+\lambda_{2}\mathcal{L}_{\operatorname{boundary}}+\mathcal{L}_{\operatorname{physics}},$ (25)
$\text{where}\quad\mathcal{L}_{\operatorname{initial}}$ $=\operatorname{RMSE}[\mathcal{G}_{\theta}(u^{b}_{0},t=0)(\boldsymbol{x}_{p})-u_{0}^{b}(\boldsymbol{x}_{p})],$
$\mathcal{L}_{\operatorname{boundary}}$ $=\operatorname{RMSE}[\mathcal{B}(\mathcal{G}_{\theta},\boldsymbol{x},t)],$
$\mathcal{L}_{\operatorname{physics}}$ $=\operatorname{RMSE}[\mathcal{R}(\mathcal{G}_{\theta}(u^{b}_{0},t)(\boldsymbol{x}_{p}),\boldsymbol{x}_{p},t)],$
where $\operatorname{RMSE}$ represents the rooted mean square error, $\mathcal{G}_{\theta}$ represents a neural operator, $\mathcal{G}$ and $\mathcal{R}$ denote the ground truth and the residual of the PDE operator, respectively. As shown in Eq. 25, $\mathcal{L}_{\operatorname{inital}}$, $\mathcal{L}_{\operatorname{boundary}}$ and $\mathcal{L}_{\operatorname{physics}}$ enforce $\mathcal{G}_{\theta}$ to satisfy the initial conditions, boundary conditions and the PDE constraints, respectively. Like PINNs [68], the PDE residuals in Eq. 25 are calculated via the auto-differentiation.
PINO [45]
PINO utilizes the pseudo-spectral or finite difference methods to construct the loss function between $\mathcal{G}_{\theta}(u^{b}_{t})$ and $\mathcal{G}_{\theta}(u^{b}_{t+\Delta t})$. PINO utilized the FNO [40] as the backbone network. The loss function in PINO can be formulated as follows:
$\mathcal{L}_{\operatorname{PINO}}$ $=\lambda\mathcal{L}_{\operatorname{boundary}}+\mathcal{L}_{\operatorname{physics}},$ (26)
$\text{where}\quad\mathcal{L}_{\operatorname{boundary}}$ $=\operatorname{RMSE}[\mathcal{B}(\mathcal{G}_{\theta},\boldsymbol{x},t)],$
$\mathcal{L}_{\operatorname{physics}}$ $=\sum_{t=0}^{T-\Delta t}\operatorname{RMSE}[\mathcal{G}_{\theta}(u^{b}_{0},t+\Delta t)(\boldsymbol{x}_{p})-\mathcal{P}(\mathcal{G}_{\theta},\boldsymbol{x}_{p},t)],$
where $\mathcal{B}$ denotes the constraints on the boundary conditions, and $\mathcal{P}$ denotes the update regime of numerical PDE solvers.
I.2: 1D Convection-Diffusion Equation
Data
The initial states $u(x,0)$ are generated from the functional space $\mathcal{F}_{N}\triangleq\{\sum_{n=1}^{N}a_{n}\sin(2\pi nx):a_{n}\sim\mathbb{U}(0,1)\},$ where $\mathbb{U}(0,1)$ denotes the uniform distribution over $(0,1)$, and $N$ represents the maximum frequency of the functional space. We generate the ground truth with the pseudo-spectral methods with the Crank–Nicolson regime. All PDE instances are generated on the spatial grid $1024$, then down-sampled to $64$. The step size is fixed as $10^{-5}$. We generate 200 test data with seed 1 and 200 validation data with seed 2.
Hyperparameters
The PINO and MCNP Solver use the 1D FNO as the backbone models. We fix the number of layers and widths as 4 and 32, respectively. We choose the best $modes$ (the number of frequency components) in $\{12,16,20\}$ for FNO, respectively. For PINO and MCNP Solver, we utilize Adam to optimize the neural network for 10000 epochs with an initial learning rate $lr$ and decay the learning rate by a factor of 0.5 every 1000 epochs. The batch size is fixed as 200. The learning rate is chosen from the set $\{0.02,0.01,0.005,0.001\}$. Because the pseudo-spectral methods can naturally stratify the boundary conditions, we fix $\lambda$ as 00 in this section. For PI-DeepONet, we choose the network structure in line with the 1D case in [90] and extend the training iterations to 50000 to ensure the convergence of the model. Moreover, we search the hyperparameters $\lambda_{1}$, $\lambda_{2}$ and the width of the neural networks in $\{1,5,10\}$, $\{1,5,10,25,50\}$ and $\{80,100,120\}$, respectively. Because PI-DeepONet-M utilizes an adaptive re-weighting scheme to balance training loss, we do not need to tune the hyper-parameters in PI-DeepONet. All hyperparameters are chosen via the validation set with seed 0.
I.3: 1D Allen-Cahn Equation
Data
The initial states $u(x,0)$ are generated from the functional space $\mathcal{F}_{N}\triangleq\{\sum_{n=1}^{N}a_{n}\sin(2\pi nx):a_{n}\sim\mathbb{U}(0,1)\},$ where $\mathbb{U}(0,1)$ denotes the uniform distribution over $(0,1)$, and $N$ represents the maximum frequency of the functional space. We generate the ground truth with the Python package ‘py-pde’ [102]. The ground truth solution is generated via the finite-difference method with the Runge-Kutta 4 method. All PDE instances are generated on the spatial grid $1024$, then down-sampled to $65$. The step size is fixed as $10^{-6}$. We generate 200 test data with seed 1 and 200 validation data with seed 2.
Hyperparameters
The PINO and MCNP Solver use the 1D FNO as the backbone models. We fix the number of layers and widths as 4 and 32, respectively. We choose the best $modes$ in $\{12,16,20\}$ for FNO, respectively. For PINO and MCNP Solver, we utilize Adam to optimize the neural network for 20000 epochs with an initial learning rate $lr$ and decay the learning rate by a factor of 0.5 every 2000 epochs. The batch size is fixed as 200. The learning rate is chosen from the set $\{0.02,0.01,0.005,0.001\}$. The hyperparameter $\lambda$ in the PINO loss is chosen from the set $\{1,2,5,10,25,50\}$. For PI-DeepONet, we choose the network structure in line with the 1D case in [90] and extend the training iterations to 200000 to ensure the convergence of the model. Moreover, we search the hyperparameters $\lambda_{1}$, $\lambda_{2}$ and the width of the neural networks in $\{1,5,10\}$, $\{1,5,10,25,50\}$ and $\{80,100,120\}$, respectively. Because PI-DeepONet-M utilizes an adaptive re-weighting scheme to balance training loss, we do not need to tune the hyper-parameters in PI-DeepONet. All hyperparameters are chosen via the validation set with seed 0.
I.4: 2D Navier-Stokes Equation
Data
We utilize the pseudo-spectral methods to generate the ground truth test data with the step size of $10^{-4}$ for the Crank–Nicolson scheme. Furthermore, all PDE instances are generated on the grid $256\times 256$, then down-sampled to $64\times 64$, which is in line with the setting in [40]. We generate 200 test data with seed 0, 200 validation data with seed 1 and 2000 training data with seed 2.
Hyperparameters
The PINO, SP-FNO and MCNP Solver use the 2D FNO as the backbone models. We fix the number of layers and widths as 4 and 36, respectively. We choose the best $modes$ in $\{12,16,20\}$ for FNO, respectively. For PINO, SP-FNO and MCNP Solver, we utilize Adam to optimize the neural network for 20000 epochs with an initial learning rate of $lr$ and decay the learning rate by a factor of 0.8 every 2000 epochs. The batch size is fixed as 10. The learning rate $lr$ is chosen from the set $\{0.02,0.01,0.005,0.001\}$. Because the pseudo-spectral methods can naturally stratify the boundary conditions, we fix the $\lambda$ as 00 in this section. For FNO, we find that a cosine annealing schedule can obtain the best result when training with the supervised regime. We utilize Adam to optimize the neural network for 400/200/100 epochs with the initial learning rate of $lr$ for FNO/FNO+/FNO++, respectively. The learning rate $lr$ is chosen from the set $\{0.02,0.01,0.005,0.001\}$. All hyperparameters are chosen via the validation set with seed 0.
I.4: 2D Fractional Diffusion Equation on a Disk
Data
The test data is generated from a fine-grained pseudo-spectral method, using the Bessel function as the basis function. The radius $[0,1]$ is divided into 400 grid points, with a time step of $10^{-4}$. Afterward, we downsample the spatiotemporal resolution to $40\times 10$. Fifty test datasets are generated using seed 1, and fifty validation datasets are generated using seed 2.
Hyperparameters
We utilize the DeepONet as a backbone model and train the MCNP Solver in a mesh-free regime. The batch size is fixed as 20. For each training epoch, we uniformly sample 2000 spatiotemporal points in the target domain and construct the loss function accordingly. We fix the number of layers and widths as 5 and 200, respectively. We use ReLU as an activation function. We utilize Adam to optimize the neural network for 10000 epochs with an initial learning rate of $lr$ and decay the learning rate by 0.8 every 2000 epochs. The learning rate $lr$ is chosen from the set $\{0.02,0.01,0.005,0.001\}$. All hyperparameters are chosen via the validation set with seed 0.
I.5: Implementation Details of Numerical Solvers
In this section, we introduce the implementation details of traditional numerical solvers, including both Eulerian and Monte Carlo methods. To ensure a fair comparison, we tested the solvers across multiple spatial resolutions, selecting the lowest resolution that produced comparable results. The implementation of traditional Eulerian methods follows the same procedure as the data generation process; however, to accelerate inference, we employed a coarser spatiotemporal grid. The specific spatiotemporal resolutions used are provided in Table VIII.
TABLE VIII: Spatiotemporal resolutions and number of particles used for testing numerical solvers.
Equation Solvers Temporal Resolution Spatial Resolution Number of Particles
Convection- Diffusion MCM / MCM+ / MCM++ 10 64 200 / 2000 / 20000
PSM / PSM+ / PSM++ 10 / 20 / 50 16
Allen- Cahn MCM / MCM+ / MCM++ 20 64 200 / 2000 / 20000
FDM / FDM+ / FDM++ 20 / 100 / 200 64
Navier- Stokes MCM / MCM+ / MCM++ 10 $64\times 64$ 10 / 20 / 50
PSM / PSM+ / PSM++ 100 / 200 / 500 $16\times 16$
Fractional MCM / MCM+ / MCM++ 20 40 200 / 2000 / 20000
PSM / PSM+ / PSM++ 10 / 20 / 50 20
Appendix  II: The Validation Loss During Training
In this section, we present the validation loss of all experiments during training process in Fig. 8.
Figure 8: The validation loss of MCNP Solver during training. CDE, ACE, and NSE are the abbreviations of the convection-diffusion equation, Allen-Cahn equation and Navier-Stokes equation, respectively. We use the results of MCNP-10, MCNP-100 and MCNP-20 to plot the training curves for CDE, ACE, and NSE, respectively.
Appendix  III: Choice of the Backbone Network
In this section, we discuss the choice of the backbone network of MCNP Solver. Firstly, we test three network structures on the 2D Navier-Stokes equation ($\nu=10^{-4}$ with Li forcing), including FNO [40], UNet [73] and MultiWaveleT- (MWT) based model [21]. We divide the time interval $[0,10]$ into ten uniform lattices for each method. Table IX presents each backbone method’s performance and computational cost on the 2D Navier-Stokes equations. In this paper, our primary contribution lies in the design of the loss function for unsupervised neural PDE-solving training. Consequently, we choose one of the most widely used backbone models, i.e., FNO, for our main experiments.
TABLE IX: Choice of the backbone network. Relative $\ell_{2}$ errors ($E_{\ell_{2}}$), relative $\ell_{\infty}$ errors ($E_{\ell_{\infty}}$) and computational costs for varying backbone models.
Model $E_{\ell_{2}}$ (%) $E_{\ell_{\infty}}$ (%) Train Time (H) Param.(M)
FNO 3.101 $\pm$0.043 5.815 $\pm$0.075 0.261 5.319
U-Net 6.030 $\pm$0.138 11.107 $\pm$0.189 0.287 6.823
MWT 3.929 $\pm$0.057 7.367 $\pm$0.086 0.621 6.361
Appendix  IV: Discussions on Hyperparameter Tuning
Compared to traditional solvers, the MCNP Solver and other neural operators require an optimization process that includes neural network training and hyperparameter tuning, while traditional numerical methods do not. We explain this issue from two perspectives. Firstly, neural operators are trained by sampling various initial fields as input, enabling them to generalize to new initial fields and predict evolving dynamics without retraining or fine-tuning after the training phase. Since neural networks can roll out quickly, neural operators can significantly accelerate the simulation process for new initial value problems. This makes them promising for inverse design and real-time physical simulations, such as weather forecasting and fluid control [83, 37]. Secondly, MCNP Solver is more efficient in hyperparameter tuning compared to other physics-driven neural operators [90, 45]. This efficiency stems from MCNP’s simple loss function, involving only Monte Carlo loss, with boundary and initial conditions naturally embedded in the random walks of the particles. For example, in models like PI-DeepONet [90], the loss includes physical, initial, and boundary terms, making hyperparameter tuning complicated. In contrast, for the MCNP Solver, the hyperparameter tuning only involves adjusting the optimizer’s learning rate and the neural operator’s $modes$ (the number of frequency components). Fig. 9 shows the performance of the Navier-Stokes equation (E3) under seed 0 with varying initial learning rates and modes, and the results show that the performance of MCNP solver with varying learning rates and modes are relatively robust, indicating easier parameter tuning.
[…]
Figure 9: The effects of hyperparameter tuning. Relative $\ell_{2}$ errors ($E_{\ell_{2}}$) for varying learning rates (a) and $modes$ (b) on Navier-Stokes experiments (E3, seed 0).
Appendix  V: Additional Simulation Results
In this section, we present a comparison between the ground truth and the predicted vorticity fields $\omega$ of all unsupervised methods from $t=2$ to $t=10$ with Zero, Li and Kolmogorov forcing, respectively (Fig. 10, Fig. 11 and 12).
Figure 10: Simulation of 2D Navier-Stokes equations. The ground truth vorticity versus the prediction of learned neural PDE solvers for an example in the test set from $t=2$ to $t=10$, with Zero forcing. (Above: $\nu=10^{-4}$; Below: $\nu=10^{-5}$).
Figure 11: Simulation of 2D Navier-Stokes equations. The ground truth vorticity versus the prediction of learned neural PDE solvers for an example in the test set from $t=2$ to $t=10$, with Li forcing. (Above: $\nu=10^{-4}$; Below: $\nu=10^{-5}$).
Figure 12: Simulation of 2D Navier-Stokes equations. The ground truth vorticity versus the prediction of learned neural PDE solvers for an example in the test set from $t=2$ to $t=10$, with Kolmogorov forcing. (Above: $\nu=10^{-4}$; Below: $\nu=10^{-5}$).